OpenStudy (anonymous):

the eye size of many vertebrates is related to the body mass by the logarithmic equation log(E) = log(10.61) + 0.1964log(m) where E is the eye axial length in millimeters, and m is the body mass in kilograms. Predict the mass of a vertebrate with an eye axial length of 22mm

OpenStudy (whpalmer4):

You want to find the mass given the formula \[\log(E) = \log(10.61) + 0.1964\log(m)\]First thing to do is rearrange that formula to give you \(m\) in terms of everything else. Can you do that?

OpenStudy (anonymous):

Not very good at this unit, so no unfortunately

myininaya (myininaya):

So E is the eye axial length (in mm) m is body mass (in kg) it wants to find m when E=22 mm

myininaya (myininaya):

Insert 22 for E and solve for m

myininaya (myininaya):

show me as far as you can get and we will work from there

OpenStudy (whpalmer4):

sorry, I must have made a mistake in my arithmetic, now it looks more reasonable.

OpenStudy (anonymous):

just double checked, and that is the formula if it helps this is a multiple choice question my options are 2.37 444.40 1.15 40.98

OpenStudy (whpalmer4):

don't be scared by the logarithms. this is just algebra. \[\log(E) = \log(10.61) + 0.1964\log(m)\]we want the term with \(m\) in it alone on one side. what would you do as a first step?

OpenStudy (whpalmer4):

How about we subtract \(\log(10.61)\) from both sides? what do you get if you do that?

OpenStudy (anonymous):

log(E) - log(10.61) = 0.1964log(m)

OpenStudy (whpalmer4):

good. now if we want to solve that for \(\log(m)\) what do we do next?

OpenStudy (anonymous):

divide both sides by 0.1964 ?

myininaya (myininaya):

yep

myininaya (myininaya):

if you want to go ahead and replace E with 22 you can an approximate the left hand side

myininaya (myininaya):

you will get something in the form \[y \approx \log(m) \implies 10^{y} \approx m\]

OpenStudy (anonymous):

alright so [log(22) - log(10.61)]/0.1974 = log(m)

myininaya (myininaya):

cool stuff and you can enter that stuff on the left hand side into a calculator i can check your approximation

OpenStudy (anonymous):

3.71306

OpenStudy (anonymous):

m = 40.9791

myininaya (myininaya):

looks like you use natural log instead not a problem we will just have to fix our base for the equivalent exponential form

myininaya (myininaya):

oh it looks like you did that

OpenStudy (whpalmer4):

Looks good to me!

myininaya (myininaya):

\[y \approx \log_e(m) \implies e^{y} \approx m\] great job